Prior to liquidating their partnership, Pepper and Morrison had capital accounts of $21,000 and $85,000, respectively. The partnership assets were sold for $40,000. The partnership had no liabilities. Pepper and Morrison share income and losses equally.

Required:

a. Determine the amount of Pepper's deficiency.
$

b. Determine the amount distributed to Morrison, assuming Pepper is unable to satisfy the deficiency.
Answer #1

Answer

  • Requirement ‘a’

--Loss on asset sold = Capital balances – Asset sold for
= 21000 + 85000 – 40000 = $ 66,000

--Income and losses are shared equally, hence, each partner will bear $ 66000 x 50% = $ 33,000

--Balances after loss distribution:
    Pepper = $ 21000 – 33000 = ($12000) [Answer to Requirement ‘a’]
   Morrison = $ 85000 – 33000 = $ 52000 [used in Requirement ‘b’]

Answer: Amount of Pepper’s Deficiency = $ 12,000

  • Requirement ‘b’

Amount distributed to Morrison = Morrison’s capital balance – Pepper’s Deficiency

= $ 52000 - $ 12000
= $ 40,000 = Answer